# Nonlinear morphoelastic energy based theory for stimuli responsive elastic shells

### v2

@@ 5.3 Eversion of a closed sphere
Inspired again by the analysis in Pezzulla et al. [Pezzulla2018PRL], we look at the effect of the same incompatible non-elastic stimulus in eq.(5.4) on a closed sphere ($s\in[-\pi,\pi]$).
@@
that allow us to recover (except for the sign) the linear behavior in $\kappa h_{0}$ highlighted in Pezzulla et al. [Pezzulla2018PRL]
This analysis allows us to reiterate the importance of retaining the geometrical non-linearities when studying the limit of small strains. In our model, the issue of retaining a bending-to-stretching coupling, recently rediscussed by Wood and Hannah [Wood2019], can be seen in Section 4.3 where we show that the coupling term identified by $w^{gnl}$ is identically zero only for intrinsically flat slender objects.
@@ 5.4 Large deformation in budding and vesiculation
Shapes associated with budding and vesiculation are typical during deformations in biological membranes and their morphologies affect their functionalities. Although the biochemical complexity behind budding in biological membranes is beyond the scope of this work, the mechanical model proposed here can be used to deepen the understanding of the role of large deformations and localized stimuli on the morphing strategies in active and biological solid-like curved shells as investigated in recent papers [Chen2025, Yu2025]. Localized control of curvature and/or area growth are indeed strategies used by biological membranes in creating protrusions that can be precursor of vesiculation [Willy2021, Zimmerberg2006]. Biological membranes are often modeled as fluid membranes [Helfrich1973] because its microscopic structure of a lipidic bi-layer allows distances to be rearranged freely and thus it cannot sustain shear loads, differently from what happens in a solid elastic membrane. However, it has been shown that elastic features must be taken into account to fully characterize the experimental evidence on the mechanical behavior of biological membranes [Deseri2013]; in those systems, the elastic behavior is thought to be due to the connection of the membranes to an underlying protein mesh [Lachowski2022, Janmey2006].
Supported by experimental data, Willy et al. [Willy2021] proposed that the formation of a vesicle from a shallow substrate is controlled by a combined in-plane growth followed by a localized curvature stimulus. Because this morphing strategy is within the class of stimuli most naturally accommodated within our modelling framework, we aim to qualitatively test what morphologies emerge in a non-linear morphoelastic spherical cap when we apply simple localized stimuli in curvature and metric, leaving a more detailed analysis to future studies.
@@
We first explore the case in which only a curvature target is imposed, i.e. we set $\gamma_{a1}=\gamma_{a2}=0$. In fig. 5(a), 5(b) and 5(c), we show the emerging morphologies that are associated with $\eta_{b}=0.5$ and $\eta_{b}=1$ for $\bar{\kappa}=300$, and $\eta_{b}=2$ for $\bar{\kappa}=128$, respectively. A target curvature that imposes a larger stimulus in the meridional direction gives rise to an elongated bulge, similar to the morphologies observed at the onset of a protrusion. An isotropic stimulus generates a more symmetric bulge, although flattened around the apex. When the target field prescribes a larger curvature along the azimuthal direction, the morphology evolves toward a ring-like bulge (in fig. 5(c) we see the onset of a trough at the apex), with a morphology reminiscent of the pits induced by clarthrin coats [Willy2021] or a caveolar superstructure [Golani2019]; imposing a larger curvature in the azimuthal direction has the effect of forcing the domain $\mathcal{I}_{s}$ to shrink towards the apex more than the meridional contribution of the stimulus can accommodate, making the formation of a ring energetically favorable over a bulge. Interestingly, the emerging morphologies display regions of both positive and negative local curvature, although the target metrics contain only positive local curvatures. Examining the elastic strains associated with the three cases, as shown in fig. 5(d), 5(e) and 5(f), we note the presence of both common and distinct features. Compressive azimuthal strains (dashed curves) emerge right outside $\mathcal{I}_{s}$ in all the cases, suggesting the potential for an instability in the hoop direction, possibly acting as a precursor to fracture [Riccobelli2024]; the meridional strain (continuous curves) remains tensile throughout the domain $\mathcal{I}_{s}$ but exhibits a sharp transition at the boundary of the region where the stimulus is applied. Within the domain $\mathcal{I}_{s}$, the strains are always tensile; however, if their maximum is always at the apex for $\eta_{b}=1$ and $\eta_{b}=2$ (with a strong localization at the apex present when the ring-like morphology), when a larger curvature is imposed along the meridional direction ($\eta_{b}=0.5$), the maximum strain is found around the middle of the meridional extent of the stimulated domain $\mathcal{I}_{s}$, potentially being the cause for the emergence of a protruded shape, instead of the more flattened apex as seen when $\eta_{b}=1$.
@@ 6 Conclusion
The formulation proposed retains geometric and constitutive nonlinearities without imposing additional kinematic constraints. In doing so, we describe the through-the-thickness behavior following the approach presented by Ozenda and Virga [Ozenda2021] and extending it to the case of a naturally curved shell. Furthermore, we embed the non-elastic morphing stimuli by decomposing them into (i) in-plane metric and curvature components and (ii) through-the-thickness component, thus enabling their independent control. The terms of order $O(h_{0}^{3})$ in the energies correctly retain all the contributions that must consistently appear at this order, including the quadratic contributions in $Z$ in the form of the functions $\phi$ and $\bar{\phi}$ as well as in the contravariant metric $\bar{\mathrm{G}}_{\alpha\beta}$.
@@ Appendix A Expansion of the invariants
In this Section we present details of results that are used, without explanation, in the main text. We recall that we consider a body $\mathcal{B}_{0}=\mathcal{S}_{0}\times[-h_{0}/2,h_{0}/2]$, where $\mathcal{S}_{0}$ defines the 2D midsurface of the object and $h_{0}$ is its thickness. Its 3D covariant metric $\overset{\circ}{\mathrm{G}}_{ij}$, with $i,j=1,2,3$, with respect to a set of a curvilinear material coordinate system $\mathbf{X}=[X^{1},X^{2},X^{3}]$, in the reference configuration takes the form
$\overset{\circ}{\mathrm{G}}_{ij}\mathrm{d}X_{i}\mathrm{d}X_{j}=\overset{\circ}{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta_{\alpha}\mathrm{d}\eta_{\beta}+Z^{2}\mathrm{d}Z\mathrm{d}Z\,$, (73)
where the metric is decoupled in the in-plane contribution, denoted by $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$, with the Greek indices $\alpha,\beta=1,2$, and the through-the-thickness contribution. We also have introduced a set of curvilinear coordinates $\eta^{1},\eta^{2},Z]$, where $\eta^{\alpha}$ (with $\alpha=1,2$) identify the in-plane coordinates, while $Z$ is the coordinate along the normal direction. As discussed in the main text, the components of the metric associated with those two contributions take the form
@@
where $\overset{\circ}{\mathrm{a}}_{\alpha\beta}$, $\overset{\circ}{\mathrm{b}}_{\alpha\beta}$, and $\overset{\circ}{\mathrm{c}}_{\alpha\beta}$ are the first, second, and third fundamental forms of the mid-surface $\mathcal{S}_{0}$, respectively. By introducing two functions $\phi(\eta^{\alpha},Z)$ and $\bar{\phi}(\eta^{\alpha},Z)$ describing the through-the-thickness behaviour in the current and intermediate configuration, respectively, we can also write the components of the covariant metrics in those two configuration as
@@
respectively. $\mathrm{a}_{\alpha\beta}$, $\mathrm{b}_{\alpha\beta}$, and $\mathrm{c}_{\alpha\beta}$ are the components of the first, second, and third fundamental forms of the mid-surface in the current configuration. $\bar{\mathrm{a}}_{\alpha\beta}$, $\bar{\mathrm{b}}_{\alpha\beta}$, and $\bar{\mathrm{c}}_{\alpha\beta}$ are the first, second, and third fundamental forms of the mid-surface in the intermediate configuration.
@@ A.1 Third invariant
In the main text we rewrite the third invariant as
@@
To derive the expansion of $\det\left(\mathrm{G}_{\alpha\beta}\right)$ and $\det\left(\bar{\mathrm{G}}_{\alpha\beta}\right)$ expanded in the functions $\phi$ and $\bar{\phi}$, we treat ${\bf G}_{\alpha\beta}$ and $\bar{\bf G}_{\alpha\beta}$ as tensorial quantities. For simplicity of notation, we show the passages on the metric $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$ expanded in the more familiar coordinate $Z$ but the same process applies for the metrics in all the configurations.
@@
$=\det(\overset{\circ}{\mathrm{a}}_{\alpha\beta})\left[1-4Z\overset{\circ}{H}+Z^{2}\left(2\overset{\circ}{K}+4\overset{\circ}{H}^{2}\right)\right]+O(Z^{3})$ (78)
@@ A.2 Second invariant
In deriving the second invariant $I_{2}$ of the right Cauchy-Green tensor ${\bf C}_{e}$
@@
where $\bar{\mathbf{g}}^{i}$ indicates the tangent vector in the intermediate configuration, with $i=1,2,3$, we want to show that
$I_{2}=\frac{1}{2}\left(\mathrm{tr}^{2}({\bf C}_{e})-\mathrm{tr}({\bf C}^{2}_{e})\right)$ (80)
@@
$I_{2}=I_{3}\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{-2}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left(I_{1}-\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\right)$ (81)
where $I_{1}$ is the first invariant of ${\bf C}_{e}$
$I_{1}$ $=\mathrm{G}_{\alpha\beta}\bar{\mathrm{G}}^{\alpha\beta}+\mathrm{G}_{33}\bar{\mathrm{G}}^{33}=\mathrm{G}_{\alpha\beta}\bar{\mathrm{G}}^{\alpha\beta}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\,.$ (82)
@@
By inserting the form of the covariant current metric in eq. (75) in eq. (79) the right Cauchy-Green tensor becomes
@@
$I_{2}=\frac{1}{2}\left(\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}-{\bf P}:\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}\right)+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)$ (89)
Applying the Cayley-Hamilton theorem to the term $\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)$, we know that
$\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)-\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)+\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right){\bf P}={\bf 0}$ (90)
and taking the (in-plane) trace with respect to $\mathbf{P}$ on both sides we obtain
${\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)-\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}+2\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)=0.$ (91)
We can thus rewrite
$\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}-{\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)=2\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right),$ (92)
and substituting this expression in eq. (89), it simplifies to
@@
$=I_{3}\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{-2}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left(I_{1}-\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\right)$ (93)
where in latter equality we used eq. (82) and eq. (83).

### v3

@@ 5.3 Eversion of a closed sphere
Inspired again by the analysis in Pezzulla et al. [25], we look at the effect of the same incompatible non-elastic stimulus in eq.(73) on a closed sphere ($s\in[-\pi,\pi]$).
@@
that allow us to recover (except for the sign) the linear behavior in $\kappa h_{0}$ highlighted in Pezzulla et al. [25]
This analysis allows us to reiterate the importance of retaining the geometrical non-linearities when studying the limit of small strains. In our model, the issue of retaining a bending-to-stretching coupling, recently rediscussed by Wood and Hanna [36], can be seen in Section 4.3 where we show that the coupling term identified by $w^{gnl}$ is identically zero only for intrinsically flat slender objects.
@@ 5.4 Large deformation in budding and vesiculation
Shapes associated with budding and vesiculation are typical during deformations in biological membranes and their morphologies affect their functionalities. Although the biochemical complexity behind budding in biological membranes is beyond the scope of this work, the mechanical model proposed here can be used to deepen the understanding of the role of large deformations and localized stimuli on the morphing strategies in active and biological solid-like curved shells as investigated in recent papers [2, 37]. Localized control of curvature and/or area growth are indeed strategies used by biological membranes in creating protrusions that can be precursor of vesiculation [35, 38]. Biological membranes are often modeled as fluid membranes [12] because its microscopic structure of a lipidic bi-layer allows distances to be rearranged freely and thus it cannot sustain shear loads, differently from what happens in a solid elastic membrane. However, it has been shown that elastic features must be taken into account to fully characterize the experimental evidence on the mechanical behavior of biological membranes [6]; in those systems, the elastic behavior is thought to be due to the connection of the membranes to an underlying protein mesh [20, 14].
Supported by experimental data, Willy et al. [35] proposed that the formation of a vesicle from a shallow substrate is controlled by a combined in-plane growth followed by a localized curvature stimulus. Because this morphing strategy is within the class of stimuli most naturally accommodated within our modelling framework, we aim to qualitatively test what morphologies emerge in a non-linear morphoelastic spherical cap when we apply simple localized stimuli in curvature and metric, leaving a more detailed analysis to future studies.
@@
We first explore the case in which only a curvature target is imposed, i.e. we set $\gamma_{a1}=\gamma_{a2}=0$. In fig. 5(a), 5(b) and 5(c), we show the emerging morphologies that are associated with $\eta_{b}=0.5$ and $\eta_{b}=1$ for $\bar{\kappa}=300$, and $\eta_{b}=2$ for $\bar{\kappa}=128$, respectively. A target curvature that imposes a larger stimulus in the meridional direction gives rise to an elongated bulge, similar to the morphologies observed at the onset of a protrusion. An isotropic stimulus generates a more symmetric bulge, although flattened around the apex. When the target field prescribes a larger curvature along the azimuthal direction, the morphology evolves toward a ring-like bulge (in fig. 5(c) we see the onset of a trough at the apex), with a morphology reminiscent of the pits induced by clarthrin coats [35] or a caveolar superstructure [9]; imposing a larger curvature in the azimuthal direction has the effect of forcing the domain $\mathcal{I}_{s}$ to shrink towards the apex more than the meridional contribution of the stimulus can accommodate, making the formation of a ring energetically favorable over a bulge. Interestingly, the emerging morphologies display regions of both positive and negative local curvature, although the target metrics contain only positive local curvatures. Examining the elastic strains associated with the three cases, as shown in fig. 5(d), 5(e) and 5(f), we note the presence of both common and distinct features. Compressive azimuthal strains (dashed curves) emerge right outside $\mathcal{I}_{s}$ in all the cases, suggesting the potential for an instability in the hoop direction, possibly acting as a precursor to fracture [26]; the meridional strain (continuous curves) remains tensile throughout the domain $\mathcal{I}_{s}$ but exhibits a sharp transition at the boundary of the region where the stimulus is applied. Within the domain $\mathcal{I}_{s}$, the strains are always tensile; however, if their maximum is always at the apex for $\eta_{b}=1$ and $\eta_{b}=2$ (with a strong localization at the apex present when the ring-like morphology), when a larger curvature is imposed along the meridional direction ($\eta_{b}=0.5$), the maximum strain is found around the middle of the meridional extent of the stimulated domain $\mathcal{I}_{s}$, potentially being the cause for the emergence of a protruded shape, instead of the more flattened apex as seen when $\eta_{b}=1$.
@@ 6 Conclusion
The formulation proposed retains geometric and constitutive nonlinearities without imposing additional kinematic constraints. In doing so, we describe the through-the-thickness behavior following the approach presented by Ozenda and Virga [24] and extending it to the case of a naturally curved shell. Furthermore, we embed the non-elastic morphing stimuli by decomposing them into (i) in-plane metric and curvature components and (ii) through-the-thickness component, thus enabling their independent control. The terms of order $O(h_{0}^{3})$ in the energies correctly retain all the contributions that must consistently appear at this order, including the quadratic contributions in $Z$ in the form of the functions $\phi$ and $\bar{\phi}$ as well as in the contravariant metric $\bar{\mathrm{G}}_{\alpha\beta}$. In the energy functionals derived here, we assume constraints on the form $\bar{\mathrm{c}}_{\alpha\beta}$ in the expansion of the metric $\bar{\mathrm{G}}_{\alpha\beta}$: those constraints can be relaxed in future work to make us able to investigate the behavior under a arbitrary intermediate metric.
@@ Appendix A Expansion of the invariants
In this Section, we present further details of results that are used in the main text. We recall that we consider a body $\mathcal{B}_{0}=\mathcal{S}_{0}\times[-h_{0}/2,h_{0}/2]$, where $\mathcal{S}_{0}$ defines the 2D midsurface of the object and $h_{0}$ is its thickness. Its 3D covariant metric $\overset{\circ}{\mathrm{G}}_{ij}$, with $i,j=1,2,3$, with respect to a set of a curvilinear material coordinate system $X^{1},X^{2},X^{3}]$, in the reference configuration takes the form
$\overset{\circ}{\mathrm{G}}_{ij}\mathrm{d}X^{i}\mathrm{d}X^{j}=\overset{\circ}{\mathrm{G}}_{\alpha\beta}\mathrm{d}\eta^{\alpha}\mathrm{d}\eta^{\beta}+Z^{2}\mathrm{d}Z\mathrm{d}Z\,$, (78)
where the metric is decoupled in the in-plane contribution, denoted by $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$, with the Greek indices $\alpha,\beta=1,2$, and the through-the-thickness contribution. We also have introduced a set of curvilinear coordinates $\eta^{1},\eta^{2},Z]$, where $\eta^{\alpha}$ (with $\alpha=1,2$) identify the in-plane coordinates, while $Z$ is the coordinate along the normal direction. As discussed in the main text, the components of the metric associated with these two contributions take the form
@@
where $\overset{\circ}{\mathrm{a}}_{\alpha\beta}$, $\overset{\circ}{\mathrm{b}}_{\alpha\beta}$, and $\overset{\circ}{\mathrm{c}}_{\alpha\beta}$ are the first, second, and third fundamental forms of the mid-surface $\mathcal{S}_{0}$, respectively. By introducing two functions $\phi(\eta^{\alpha},Z)$ and $\bar{\phi}(\eta^{\alpha},Z)$ describing the through-the-thickness behavior in the current and intermediate configuration, respectively, we can also write the components of the covariant metrics in these two configurations as
@@
We notice that $\mathrm{a}_{\alpha\beta}$, $\mathrm{b}_{\alpha\beta}$, and $\mathrm{c}_{\alpha\beta}$ are the components of the first, second, and third fundamental forms of the mid-surface in the current configuration, while $\bar{\mathrm{a}}_{\alpha\beta}$, $\bar{\mathrm{b}}_{\alpha\beta}$, and $\bar{\mathrm{c}}_{\alpha\beta}$ are the components of the target first, second, and third forms of the mid-surface in the intermediate configuration.
@@ A.1 Third invariant
In the main text, we rewrite the third invariant as
@@
To derive the expansions of $\det\left(\mathrm{G}_{\alpha\beta}\right)$ and $\det\left(\bar{\mathrm{G}}_{\alpha\beta}\right)$ in the functions $\phi$ and $\bar{\phi}$, we treat ${\bf G}_{\alpha\beta}$ and $\bar{\bf G}_{\alpha\beta}$ as tensorial quantities. For simplicity of notation, we show the passages on the metric $\overset{\circ}{\mathrm{G}}_{\alpha\beta}$ expanded in the more familiar coordinate $Z$ but the same process applies for the metrics in all configurations.
@@
$=\det(\overset{\circ}{\mathrm{a}}_{\alpha\beta})\left[1-4Z\overset{\circ}{H}+Z^{2}\left(2\overset{\circ}{K}+4\overset{\circ}{H}^{2}\right)\right]+O(Z^{3})\,$, (83)
@@ A.2 Second invariant
In deriving the second invariant $I_{2}$ of the right Cauchy-Green tensor ${\bf C}_{e}$, written as
@@
where $\bar{\mathbf{g}}^{i}$ indicates the tangent vector in the intermediate configuration, we want to show that
$I_{2}=\frac{1}{2}\left(\mathrm{tr}^{2}({\bf C}_{e})-\mathrm{tr}({\bf C}^{2}_{e})\right)\,$, (85)
@@
$I_{2}=I_{3}\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{-2}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left(I_{1}-\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\right)\,$, (86)
where $I_{1}$ is the first invariant of ${\bf C}_{e}$, expressed as
$I_{1}$ $=\mathrm{G}_{\alpha\beta}\bar{\mathrm{G}}^{\alpha\beta}+\mathrm{G}_{33}\bar{\mathrm{G}}^{33}=\mathrm{G}_{\alpha\beta}\bar{\mathrm{G}}^{\alpha\beta}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\,$, (87)
@@
By inserting the form of the covariant current metric, as expressed in eq. (17), in eq. (84), the right Cauchy-Green tensor becomes
@@
$I_{2}=\frac{1}{2}\left(\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}-{\bf P}:\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}\right)+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)\,.$ (94)
Applying the Cayley-Hamilton theorem to the term $\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)$, we can write
$\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)-\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)+\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right){\bf P}={\bf 0}\,.$ (95)
Moreover, taking the (in-plane) trace with respect to $\mathbf{P}$ on both sides, we obtain
${\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)-\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}+2\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)=0\,$, (96)
which we can thus rewrite as
$\left({\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)^{2}-{\bf P}:\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)=2\mathrm{det}\left(\mathrm{G}_{\alpha\beta}\left(\bar{\mathbf{g}}^{\alpha}\otimes\bar{\mathbf{g}}^{\beta}\right)\right)\,.$ (97)
Finally, substituting this expression in eq. (94), the second invariant can be written as
@@
$=I_{3}\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{-2}+\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\left(I_{1}-\left(\frac{\phi^{\prime}}{\bar{\phi}^{\prime}}\right)^{2}\right)\,$, (98)
where, in the latter equality, we used eq. (87) and eq. (88).